7. Untied and Air R Us are the only two airlines operating flights between Collegeville and Bigtown. That is, they operate
Answer #1

a).

Consider the given problem here there are two players “P1=Untied” and P2= Air R Us”. So, of then have two possible strategies “LP=low price” and “HP=high price”. Now, if “P1” charge LP then the optimum price for “P2” is “LP” because the “P2” gets “$20” as profit which is more than “$0”. Similarly, if “P1” charge HP then the optimum price for “P2” is “LP” because the “P2” gets “$50” as profit which is more than “$40”. So, in both the cases “P2” choose “LP”, => “LP” is the dominant strategy for “P2”. Similarly with the same logic we can prove that “LP” is the dominant strategy for “P1”, => the only NE is “LP, LP”, => both the player will end uo by choosing “LP” and will get “$20” as a profit.

b).

i).

Now, let’s assume that the same game will be played twice and “Untied=P1” always charge LP when “P2 = Air R Us” also play LP, => in both the round both the player will choose “LP” and will get “$20” as a profit, => the payoff are given by “20+20=$40”. So, at the end of 2 round both the player will get “$40” as a profit.

ii).

Now, “P1” always choose “LP” but “P2” choose “tit for tat”, => in the 1st round “P1” choose “LP” and “P2” choose “HP”, => “P1” gets “$50” and “P2” gets “$0” as a profit. Now, in the 2nd round “P1” choose “LP” and “P2” choose “LP”, => both the players gets “$20” as a profit. So, the payoff under this case is “50+20=$70” for “P1” and “$20” for “P2”.

iii).

Now, “P1” always choose “tit for tat” but “P2” choose “LP”, => in the 1st round “P1” choose “HP” and “P2” choose “LP”, => “P1” gets “$0” and “P2” gets “$50” as a profit. Now, in the 2nd round “P1” choose “LP” and “P2” choose “LP”, => both the players gets “$20” as a profit. So, the payoff under this case is “0+20=$20” for “P1” and “$50+$20=$70” for “P2”.

iv).

Now, if both the players choose “tit for tat”, => in the 1st round both choose “HP”, => both gets “$40” as a profit. Now, in the 2nd round both further choose “HP”, => both the player gets “$40” as a profit in the 2nd round. So, the payoff under this case is “40+40=$80”, => both the player gets “$80” as a profit at the end of the 2nd round.